The amount of time devoted to preparing for a statistics examination by students is a normally distributed random variable with a mean of 17 hours and a standard deviation of 5 hours.


Required:

a) What is the amount of time below which only 15% of all students spend studying?

b) What is the amount of time above which only one third of all students spend studying? c) What is the probability that a student spends between 16 and 20 hours studying?

d) What is the probability that a student spends at least 15 hours studying?

e) What is the probability that a student spends at most 18 hours studying?


1
Expert's answer
2022-05-12T04:06:50-0400

a)


"P(Z<\dfrac{x-17}{5})=0.15"

"\dfrac{x-17}{5}\approx-1.0364"

"x\approx11.818"

b)


"P(Z<\dfrac{x-17}{5})=1/3"

"\dfrac{x-17}{5}\approx-0.4307"

"x\approx14.8465"

c)


"P(16<X<20)=P(Z<\dfrac{20-17}{5})"

"=P(Z<0.6)-P(Z\le-0.2)"

"\approx0.72575-0.42074\approx0.3050"

d)


"P(X\ge15)=1-P(Z<\dfrac{15-17}{5})"

"=1-P(Z<-0.4)"

"\approx0.6554"



e)


"P(X\le18)=P(Z\le\dfrac{18-17}{5})"

"=P(Z\le0.2)\approx0.5793"